# See What You Seek: Semantic Contextual Integration for Cloth-Changing Person Re-Identification

### v1

@@ V Conclusion and Limitation
In this study, we propose the Semantic Contextual Integration (SCI) network, a novel method incorporates prompt learning to address cloth-changing person re-identification (CC-ReID). The SCI network includes a Semantic Separation Enhancement (SSE) module that generates ID-specific and clothing prompts to effectively isolate clothing semantics (negative factors) from ID information, thereby preserving high-dimensional semantic details (positive factors) embedded in visual images. Additionally, the Semantic-Guided Interaction Module (SIM) leverages textual cues to guide and enhance the discrimination of visual features, facilitating seamless interaction between visual and textual branches. We believe this work provides valuable insights and potential directions for future research into the role of prompt learning in advancing CC-ReID.
While learnable prompts enable the model to autonomously construct and learn, have indeed demonstrated superior outcomes and a significant enhancement in performance compared to the manually crafted prompts used in CLIP, the prompts derived by the model may lack semantic clarity and remain illegible to human observers. As a result, it is not feasible to visually represent the learned content in a textual format. Thus, future research will need to address the interpretability of these prompts and investigate methodologies that enhance their comprehensibility while maintaining their efficacy.

### v2

@@ V Conclusion and Limitation
In this study, we propose the Semantic Contextual Integration (SCI) network, a novel method that incorporates prompt learning to address cloth-changing person re-identification (CC-ReID). The SCI network includes a Semantic Separation Enhancement (SSE) module that generates ID-specific and clothing prompts to effectively isolate clothing semantics (negative factors) from ID information, thereby preserving high-dimensional ID cues (positive factors) in visual features. Additionally, the Semantic-Guided Interaction Module (SIM) leverages these textual cues to guide and enhance the discrimination of visual features, facilitating seamless interaction between the visual and textual branches. Our work provides valuable insights and suggests promising directions for future research on the role of prompt learning in CC-ReID.
Although learnable prompts outperform the manually crafted prompts used in CLIP, they often lack semantic clarity and remain uninterpretable to human observers. Consequently, the learned prompts cannot be directly visualized in a human-readable format. Future research should address the interpretability of these prompts and develop techniques to enhance their comprehensibility without compromising their discriminative power.